OpenStudy (anonymous):

express x in terms of e: ln x^2 =10

OpenStudy (amistre64):

2 = e^10....

OpenStudy (amistre64):

if you know your log and exponent rules, its just that simple

OpenStudy (anonymous):

okay thank you

OpenStudy (amistre64):

i didnt read the question good enough....hold on :)

OpenStudy (amistre64):

ln(x^2) = 10 2 ln(x) = 10 ln(x) = 10/2 ln(x) = 5 e^ln(x) = e^5 x = e^5

OpenStudy (anonymous):

awesome! stepps help hugee !!

OpenStudy (amistre64):

so do glasses :)

OpenStudy (anonymous):

:)